OpenStudy (anonymous):

When you have the basic parametric equation of: v(t)=x(t)i+y(t)j Then does r'(t)=x'(t)i+y'(t)j or x(t)i+y(t)j?

OpenStudy (anonymous):

My guess is that \(r'(t) = v(t) = x(t)\hat{\imath}+y(t)\hat{\jmath} \)

OpenStudy (anonymous):

But you definitely need more information.

OpenStudy (anonymous):

Assuming \(r(t)\) is the position and \(v(t)\) is the velocity, then \(r'(t) = v(t)\).

OpenStudy (anonymous):

Ohhhhh. Thanks. I just realized that they are implying that v is the velocity. Thank you.